A particle starts from rest with uniform acceleration a. Its velocity after n seconds is v. the displacement of the particle in the last two seconds is.....

Grade:11

6 Answers

erra akhil
8 years ago
Dear Pranjal,
v = u + a t -----eqn.1
v-final velocity
u-initial velocity
a-acceleration
t-time taken
let the velocity of the particle be v` just two seconds before which means t=(n-2) seconds
Since initial velocity=0.
on applying eq.1,
v`=0+a(n-2)
Now,
S = u t + (1/2) a t2
So the displacement in the last two seconds is
here the velocity before two seconds as found above is {(n-2)a} which becomes initial velocity for last 2 seconds and time is 2 seconds because we want displacement for last 2 seconds
S = {a (n-2) } 2 + (½) a (22)
S = 2 a (n-2) + 2a
S= 2an-2a
S= 2a(n-1) for last two seconds.
 
erra akhil
8 years ago
Approve if my answer helped you
 
Srinivas
5 years ago
A particle starts from rest moving with uniform acceleration and covers 26m in 7th Second. Find it's displacement 9th second
Raushan kumar
5 years ago
A particle is moving with constant acceleration from A to B in a straight line AB .If u and v are the velocities at A and B respectively, then it's velocity at the midpoint c will be 
Raushan kumar
5 years ago
A particle is moving with constant acceleration from A to B in a straight line AB. If u and v are the velocities at A and B respectively , then it's velocity at the midpoint c will be
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your question.
 
v = u + a t -----eqn.1
v-final velocity
u-initial velocity
a-acceleration
t-time taken
let the velocity of the particle be v` just two seconds before which means t=(n-2) seconds
Since initial velocity=0.
on applying eq.1,
v`=0+a(n-2)
Now,
S = u t + (1/2) a t2
So the displacement in the last two seconds is
here the velocity before two seconds as found above is {(n-2)a} which becomes initial velocity for last 2 seconds and time is 2 seconds because we want displacement for last 2 seconds
S = {a (n-2) } 2 + (½) a (22)
S = 2 a (n-2) + 2a
S= 2an-2a
S= 2a(n-1) for last two seconds.
 
Thanks and regards,
Kushagra